OpenStudy (anonymous):

Prove that : sin⁡3θ/sin⁡θ - cos⁡3θ/cos⁡θ =2

OpenStudy (ahsome):

Is this the equation? \[\frac{sin*3*\theta}{sin*\theta}-\frac{cos*3*\theta}{cos*\theta}=2\]

OpenStudy (loser66):

From the left hand side 1) Make them have the same denominator 2/ Use sin (a-b) = sin a cos b- sin b cos a to manipulate the numerator 3/ Use sin 2a = 2 sina cos a to manipulate the denominator 4/ Cancel out and get the right hand side. Work out and show your attempt, please.

OpenStudy (anonymous):

i know the best way take a nap and refresh yourself then u will be able to study good

OpenStudy (shamim):

sin 3 theta=?